Give an example of a function f: Σ* x N -> Z , where Σ = {a,b} and where N represents the non-negative integers and Z represents the integers.
Answer #1

There are multiple such function. Consider the input of the form string w \in \Sigma^* and non negative integer N, and f outputs |w|-N i.e. the difference of size of input string w with non negative integer N. Clearly the output will be integer.

Ex:- f(abba,2) = 4-2 =2 and f(ababab,10)=6-10= -4

Please comment for any clarification